OpenStudy (anonymous):

write the standard form equation for the line through (5,1) parallel to y=(3/4)x-7

OpenStudy (anonymous):

parallel to the line means the same slope, right? slope of your given line (number in a front of x) = 3/4 so, your line through (5,1) should have the same slope, so: y-1=3/4 *(x-5) y=3x/4 -15/4 +1 y=(3/4)x - 11/4